# Fish-Speech: Leveraging Large Language Models for Advanced Multilingual Text-to-Speech Synthesis

### v1

@@ Abstract
Keywords Text to Speecch  $\cdot$ LLM  $\cdot$ Voice Cloning
@@ 1 Introduction
The past decade has seen remarkable progress in Text-to-Speech (TTS) systems, transforming applications from virtual assistants to educational tools. Current TTS architectures, such as VALL-E [Wang et al. (2023)], VITS [Kim et al. (2021)], Fastspeech [Ren et al. (2020)] typically rely on grapheme-to-phoneme (G2P) conversion [Klatt (1987)] to convert text into phonetic representations before synthesis. While effective, this approach struggles with context-dependent polyphonic words and cross-lingual generalization due to complex phonetic rules. Recent advances in zero-shot voice conversion such as YourTTS [Casanova et al. (2022)] and unified speech generation model UniAudio [Yang et al. (2023)] have shown the potential of neural architectures in handling various speech tasks. Additionally, flow-based models like CosyVoice[Du et al. (2024), MatchaTTSMehta et al. (2024)] have demonstrated promising results in natural speech synthesis. However, most solutions disentangled semantic and accoustic feature as a trade-of to improve stability, and reduce the voice cloning / bg understanding ability.
@@ 3.2 Firefly-GAN
Firefly-GAN (FF-GAN) is an enhanced version of the EVA-GAN architecture with significant structural improvements. The model implements an optimized structure that replace the conventional convolutional and transposed convolutional components found in HiFi-GAN [Kong et al. (2020)]. We use ParallelBlock to subtitute for Multi-Receptive Field (MRF), which specifically engineered for typo-codebook vocoder applications. Through the integration of the Grouped Finite Scalar Vector Quantization (GFSQ) methodology, FF-GAN demonstrates enhanced sequence generation stability, effectively improving complex linguistic variations and facilitating natural multilingual synthesis. This architectural innovation distinguishes FF-GAN from traditional vocoders, particularly in text-to-speech synthesis and voice cloning applications within AI agent frameworks.
@@ 3.2.1 Firefly Generator
In our architecture, the conventional Multi-Receptive Field (MRF) module is replaced by a ParallelBlock, optimizing typo-codebook input processing efficiency. The ParallelBlock implements configurable convolution kernel sizes and dilation rates Yu (2015), utilizing stack-and-average mechanism for processing outputs from three ResBlocks, instead of directly addition operations. ParallelBlock Liao et al. (2024) offers enhanced receptive field coverage, superior feature extraction capabilities, and improved configurability, contributing to higher-quality audio synthesis.
@@ 3.2.2 Quantization Techniques
In order to fit the typo-codebook task, we use Grouped Finite Scalar Vector Quantization (GFSQ) to act as a vq codebook in our system. The text below will elaboratly tell you how we develop the GFSQ.
@@ 3.2.3 Conclusion
Our implementation of using GFSQ techniques achieves neaely 100% codebook utilization, and gain better objetive and subjective score in our inner ablation than other quantization techiques like RFSQ, RVQ and GRFSQ. FF-GAN significantly enhancing stability in typo-codebook operations and ensuring comprehensive intermediate variable information retention in multi-emotional and multilingual tasks.
FF-GAN’s innovative approach to typo-codebook stability has being already used in various song and music generation applications. The framework’s performance and architectural could make it position it as a reference model for future AI agent development.
@@ 4.2 Inference
Our inference strategy follows the architecture in Fig. 1. Using fish-tech including KV-cache [Pope et al. (2023)], torch compile and other acceleration methodologies, the system achieves real-time factors of approximately 1:5 on consumer-grade NVIDIA RTX 4060 mobile platforms and 1:15 on high-performance NVIDIA RTX 4090 configurations. These architectural optimizations significantly improve latency in infering, achieving a first-packet latency of 150ms.
@@ 6 Experimental Evaluation
 [^1] For experimental validation, we constrained our analysis to monolingual voice cloning scenarios, excluding cross-lingual synthesis tasks. The evaluation corpus comprised 10 distinct speaker (include different languages) identities, with 30 synthesized utterances per speaker, yielding a comprehensive evaluation set of 300 samples. It should be noted that cross-linguistic synthesis was not included in this evaluation.
@@ 6.1 Word Error Rate Analysis
Model Name WER(%)
@@
11footnotetext: Results obtained using OpenAI Whisper-medium ASR model for transcription evaluation.
@@ 6.2 Speaker Similarity Analysis
Model Name Resemblyzer SpeechBarin [^2] Experiments were conducted using the SpeechBrain framework Ravanelli et al. (2024), version 1.0.1, as the underlying speech processing toolkit.
@@
Table 2 shows the effect of our typo-codebook strategy on speaker similarity metrics. Our fish-speech model achieves similarity scores of 0.914 and 0.762 on Resemblyzer and SpeechBrain, respectively, which are remarkably close to the ground truth performance (0.921 and 0.770). The gap of only 0.76% from the ground truth in Resemblyzer and 1.04% in SpeechBrain evaluations shows the superior ability of our model to capture natural speech characteristics. The results strongly suggest that our typo-codebook architecture enables a more comprehensive capture of acoustic states, leading to improved timbral fidelity of synthesized speech. Our approach significantly outperforms baseline models such as F5-TTS (0.905 and 0.787) and rechoo (0.887 and 0.636). The consistent performance in both evaluation frameworks proves the validity of our method in preserving speaker features, which is crucial for high-quality text-to-speech synthesis and agent tasks.
@@ 8 Future Work
Building on these foundations, we propose several directions for future research. We plan to improve the performance of our model by integrating reinforcement learning techniques, focusing on improving cross-lingual generalization and emotional stability. We are also developing the Fish Agent application, an end-to-end language model based on our Fish-Speech framework. A preliminary demonstration of this system is currently available at fish.audio/demo/live. We remain committed to the open source community and will continue to maintain and extend our codebase to provide broader access to these technologies for researchers and developers.
@@ Appendix A Training Details
We trained our model on an NVIDIA H100 GPU with the following hyperparameters:

### v2

@@ Abstract
Keywords Text to Speech  $\cdot$ LLM  $\cdot$ Voice Cloning
@@ 1 Introduction
The past decade has seen remarkable progress in Text-to-Speech (TTS) systems, transforming applications from virtual assistants to educational tools. Current TTS architectures, such as VALL-E [Wang et al. (2023)], VITS [Kim et al. (2021)], FastSpeech [Ren et al. (2020)] typically rely on grapheme-to-phoneme (G2P) conversion [Klatt (1987)] to convert text into phonetic representations before synthesis. While effective, this approach struggles with context-dependent polyphonic words and cross-lingual generalization due to complex phonetic rules. Recent advances in zero-shot voice conversion such as YourTTS [Casanova et al. (2022)] and unified speech generation model UniAudio [Yang et al. (2023)] have shown the potential of neural architectures in handling various speech tasks. Additionally, flow-based models like CosyVoice[Du et al. (2024), MatchaTTSMehta et al. (2024)] have demonstrated promising results in natural speech synthesis. However, most solutions disentangled semantic and acoustic feature as a trade-off to improve stability, and reduce the voice cloning understanding ability.
@@ 3.2.1 Firefly Generator
In our architecture, the conventional Multi-Receptive Field (MRF) module is replaced by a ParallelBlock, optimizing typo-codebook input processing efficiency. The ParallelBlock implements configurable convolution kernel sizes and dilation rates, utilizing stack-and-average mechanism for processing outputs from three ResBlocks, instead of directly addition operations. ParallelBlock Liao et al. (2024) offers enhanced receptive field coverage, superior feature extraction capabilities, and improved configurability, contributing to higher-quality audio synthesis.
@@ 3.2.2 Quantization Techniques
In order to fit the typo-codebook task, we use Grouped Finite Scalar Vector Quantization (GFSQ) to act as a vq codebook in our system. The text below will elaborately tell you how we develop the GFSQ.
@@ 3.2.3 Conclusion
Our implementation of using GFSQ techniques achieves nearly 100% codebook utilization, and gain better objetive and subjective score in our inner ablation than other quantization techiques like RFSQ, RVQ and GRFSQ. FF-GAN significantly enhancing stability in typo-codebook operations and ensuring comprehensive intermediate variable information retention in multi-emotional and multilingual tasks.
FF-GAN’s innovative approach to typo-codebook stability has being already used in various song and music generation applications. The framework’s performance and architectural could make it as a reference model for future AI agent development.
@@ 4.2 Inference
Our inference strategy follows the architecture in Fig. 1. Using fish-tech including KV-cache [Pope et al. (2023)], torch compile and other acceleration methodologies, the system achieves real-time factors of approximately 1:5 on consumer-grade NVIDIA RTX 4060 mobile platforms and 1:15 on high-performance NVIDIA RTX 4090 configurations. These architectural optimizations significantly improve latency in inferring, achieving a first-packet latency of 150ms.
@@ 6 Experimental Evaluation
 [^1] For experimental validation, we constrained our analysis to monolingual voice cloning scenarios, excluding cross-lingual synthesis tasks. The evaluation corpus comprised 10 distinct speakers (include different languages) identities, with 30 synthesized utterances per speaker, yielding a comprehensive evaluation set of 300 samples. It should be noted that cross-linguistic synthesis was not included in this evaluation.
@@ 6.1 Word Error Rate Analysis
Model Name WER(%)  [^2] Results obtained using OpenAI Whisper-medium ASR model for transcription evaluation.
@@ 6.2 Speaker Similarity Analysis
Model Name Resemblyzer SpeechBrain [^3] Experiments were conducted using the SpeechBrain framework Ravanelli et al. (2024), version 1.0.1, as the underlying speech processing toolkit.
@@
Table 2 shows the effect of our typo-codebook strategy on speaker similarity metrics. Our fish-speech model achieves similarity scores of 0.914 and 0.762 on Resemblyzer and SpeechBrain, respectively, which are remarkably close to the ground truth performance (0.921 and 0.770). The gap of only 0.76% from the ground truth in Resemblyzer and 1.04% in SpeechBrain evaluations shows the superior ability of our model to capture natural speech characteristics. The results strongly suggest that our typo-codebook architecture enables a more comprehensive capture of acoustic states, leading to improved timbral fidelity of synthesized speech. Our approach significantly outperforms baseline models such as F5-TTS (0.905 and 0.787) and reecho (0.887 and 0.636). The consistent performance in both evaluation frameworks proves the validity of our method in preserving speaker features, which is crucial for high-quality text-to-speech synthesis and agent tasks.
@@ 8 Future Work
Building on these foundations, we propose several directions for future research. We plan to improve the performance of our model by integrating reinforcement learning techniques, focusing on improving cross-lingual generalization and emotional stability. We are also developing the Fish Agent application, an end-to-end language model based on our Fish-Speech framework. A preliminary demonstration of this system is currently available at fish.audio/demo/live. We remain committed to the open source community and will continue to maintain and extend our codebase to provide broader access to these technologies for researchers and developers.br
@@ Appendix A Training Details
We trained our model on NVIDIA H100 cluster with the following hyperparameters: